### v1

@@ Abstract
Replicator dynamics have widely been used in evolutionary game theory to model how strategy frequencies evolve over time in large populations. The so-called payoff matrix encodes the pairwise fitness that each strategy obtains when interacting with every other strategy, and it solely determines the replicator dynamics. If the payoff matrix is unknown, we show in this paper that it cannot be inferred from observed strategy frequencies alone — distinct payoff matrices can induce the same replicator dynamics. We thus look for a canonical representative of the payoff matrix in the equivalence class. The main result of the paper is to show that for every polynomial replicator dynamics (i.e., the vector field is a polynomial), there always exists a skew-symmetric, polynomial payoff matrix that can induce the given dynamics.
@@ 1 Introduction
The so-called replicator dynamics, originally introduced in [7], have widely been used in evolutionary game theory [9, 20] to describe how strategy frequencies change over time in a large population. Consider the basic setting where every agent in the population has $n$ (common) pure strategies at its disposal at any time. Denote by $x_{i}(t)$ the proportion of the agents taking the $i$th strategy at time $t$. Let $x(t):=(x_{1}(t),\ldots,x_{n}(t))$ be the system state. Note that $\sum_{i=1}^{n}x_{i}(t)=1$ and $x_{i}(t)\geq 0$ for all $i=1,\ldots,n$, so $x(t)$ is a probability vector. The agents in the population randomly pair with each other to play symmetric games whose payoff matrix is given by $H(x(t))\in\mathbb{R}^{n\times n}$ (i.e., the payoff matrix is a function of the current state $x(t)$). The payoff matrix encodes the competitive/cooperative nature between different strategies. Specifically, the $ij$th entry of $H(x(t))$ represents the payoff received by any agent using strategy $i$ when it interacts with an opponent using strategy $j$ with $x(t)$ the population state.
When playing the symmetric game, each agent evaluates the incremental benefit of abandoning its current strategy $i$ in favor of an alternative $k$. Conditional on facing an opponent who plays $j$, this benefit is the payoff difference $H_{kj}(x(t))-H_{ij}(x(t))$. Note that the strategy revision depends exclusively on such differences. Thus, it is invariant to any column-wise uniform shift of the payoff matrix (i.e., only the relative values of the entries of the $j$th column of $H(x(t))$ matter). Agents then follow a prescribed learning rule (e.g., pairwise proportional comparison and imitation driven by dissatisfaction [17]) to revise their strategies. Taking an appropriate limit (see, e.g., [10] for detail), we arrive at the following ordinary differential equation for $x(t)$:
@@
where $p(x)=(p_{1}(x),\ldots,p_{n}(x)):=H(x)x$ is known as the payoff function [14] and $\bar{p}(x):=\sum_{j=1}^{n}x_{j}p_{j}(x)$ is the average payoff. The dynamics of $x(t)$ evolve on the standard simplex $\Delta^{n-1}:=\{x\in\mathbb{R}^{n}\mid x_{i}\geq 0\mbox{ and }\sum_{i=1}^{n}x_{i}=1\}$.
@@
In this paper, we take a different perspective: Assuming that one can only observe the dynamics of $x(t)$, we investigate whether it is feasible to identify the unknown payoff matrix $H:\Delta^{n-1}\to\mathbb{R}^{n\times n}$? Said in other words, we ask for any given replicator dynamics, is there a unique payoff matrix associated with it? If the answer is negative, then is there a “canonical” representative (we will make the canonicity clear)? The problems we pose play important roles in the setting where a monitor in their study of ecosystem aims to uncover the competitive/cooperative relations between different species, or, the setting where a central planner in the congestion game aims to understand the agents’ incentives against others’ choice of strategies. We prove that for any replicator dynamics, there exists a corresponding zero-sum population game representation. Specifically, we show that for a generalized class of replicator systems $\dot{x}=\operatorname{diag}(x)\,g(x),x\in\Delta^{n-1},$ where $g:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a polynomial vector field satisfying $x^{\!\top}g(x)=0$ for all $x\in\Delta^{n-1}$, there exists a skew-symmetric matrix $A(x)$ such that $g(x)=A(x)\,x,\text{with }A(x)=-A(x)^{\!\top}.$ This representation produces an equivalent zero-sum population game where the payoff for each pair of agents is skew-symmetric, and the average payoff of the population is identically zero.
The zero-sum population games, defined by skew-symmetric payoff matrices, capture pure competition in which every gain for one strategy is matched by an equal loss to the rest. This structure yields a benchmark devoted to competitive dynamics while remaining tractable for the analysis of stability, coexistence [11], and learning behavior [12]. For constant payoffs, the canonical case is the three-strategy Rock-Paper-Scissors(RPS) games and its $n$-strategy generalizations. The associated replicator dynamics exhibit recurrent interior orbits on the simplex and constitute the standard framework in which population permanence [1, 8] has been rigorously established.
@@
The replicator dynamics has been studied in various scenarios. In [11, 19], the authors investigated the co-evolution of the replicator system where the payoff matrix is generated from a dynamic system based on a generalized RPS game. The natural transformation [9] was performed to study the cyclic behavior. In [14, 13], the authors studied a more general setting where the payoff function is generated from a dynamic system, and the passivity theorem [2, 16] is used to find the equilibria of the system. The most related work to ours is the static payoff model [15], where the payoff is a static function. For $p(x)$ being static, the dynamic behavior of (1) has been extensively investigated in the literature. Many existing works have focused on the case where the payoff matrix $H(x)$ is constant: For $H$ a symmetric matrix, it has been shown [17] that the replicator dynamics (1) are gradient flows of the potential function $\bar{p}(x)$ with respect to the Shahshahani metric. For $H$ skew-symmetric, it has been shown [4] that there is a unique global attractor of (1) (roughly speaking, an attractor is a minimal compact set that is forward invariant and asymptotically stable). For the case where $H$ is a 33-by-33 skew-symmetric matrix, a complete characterization of the attractors has been obtained by Zeeman [22]. In fact, more has been shown in [22] as well as in the follow-up work [5, 6] for $H$ a general 33-by-33 matrix. Amongst others, they have presented, as a conjecture with partial proofs, all possible phase portraits (49 in total) for the resulting replicator dynamics with isolated equilibria and periodic orbits.
@@
The literature is relatively sparse for the case where the payoff matrix $H(x)$ depends on the state $x$, which is more realistic for a number of settings. For example, such dependence arises naturally in modeling the ecosystem where the interactions among different species are often in higher order [3] (in this context, the $x_{i}(t)$’s represent the normalized population densities of the species). Also, for congestion games [17] where different strategies represent different routes between the destinations, it is easy to see that the payoff matrix $H(x)$ depends on $x$ (the $x_{i}$’s represent the strategy frequencies). We further mention the demand-response program for smart grids [14, Example 2] where the payoff matrix $H(x)$ depends on the user-strategy frequencies. The analysis of replicator dynamics for non-constant $H(x)$ is, however, much more involved and can quickly lose tractability. Results have been obtained for some special 33-by-33 payoff matrices $H(x)$ in [21] where $H(x)$ is an affine function. Specifically, the authors have exhibited all the equilibria of the resulting replicator dynamics, analyzed their local stability by evaluating the Jacobian matrices, and carried out numerical studies for the phase portrait.
@@
We introduce the notation that will be used throughout the analysis. We define the hyperplane $\mathbf{H}:=\{x\in\mathbb{R}^{n}|\sum_{i=1}^{n}x_{i}=1\}$. Given arbitrary integer $d>0$, we let $\mathbf{\Omega}_{[d]}^{k\times k}$ be the vector space of all $k\times k$ skew-symmetric matrices $A$ with $A^{\top}=-A$, whose entries are polynomials in $x\in\mathbb{R}^{n}$ of degree at most $d$, and $\boldsymbol{\Omega}_{(d)}^{k\times k}$ denote the homogeneous subspace of degree $d$. If $d=0$, we write $\mathbf{\Omega}^{n\times n}$ for simplicity.
@@
${\mathbf{\Omega}^{k\times k}_{[d]}=\oplus_{m=1}^{d}\mathbf{\Omega}^{k\times k}_{(m)}\ \mbox{ and }\ }\mathbf{P}^{k\times\ell}_{[d]}=\oplus_{m=1}^{d}\mathbf{P}^{k\times\ell}_{(m)}.$
For the case where $d<0$, we set $(d)$ as $(0)$.
We use multi-index $x^{\alpha}$ to denote the monomial $x^{\alpha}:=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}$, where $\alpha=\left(\alpha_{1},\ldots,\alpha_{n}\right)\in\mathbb{N}^{n}$ and $|\alpha|:=\sum_{i=1}^{n}\alpha_{i}$ is the degree of the monomial. We say that a scalar polynomial contains a monomial if the coefficient of the monomial in the standard multivariate expansion of the polynomial is non-zero. The degree of the polynomial is by convention the largest degree of a monomial contained in it. For a vector- or matrix-valued polynomial, we define its degree to be the largest degree among all of its entries.
2 Main Result
@@ 2.1 Motivating Example
$\phi:H(x)\mapsto f(x):=\operatorname{diag}(x(t))(H(x)x-x^{\top}H(x)x\mathbf{1}),$ (2)
@@ 2.2 Main Result
then, there exists a polynomial map $A:\mathbb{R}^{n}\to\mathbf{\Omega}^{n\times n}$ of degree at most $d$ such that $g(x)=A(x)x$.
@@
To wit, $x^{\top}g(x)$ is a polynomial function (and hence, analytic) and it is identically zero on $\Delta^{n-1}$, whose interior is open in $\mathbf{H}$. It then follows from the identity theorem [18] that $x^{\top}g(x)=0$ for all $x\in\mathbf{H}$.
@@
It is clear that all such $g(x)$ of degree $(d+1)$ form a vector space, which we denote by $\mathbf{Q}_{(d+1)}^{n}$. We will provide a complete characterization of the vector space by exhibiting a basis of it. In particular, we will show in Lemma 3.5 that each basis element takes the following form:
@@
where $\alpha$ is some multi-index with $|\alpha|=d$. It then follows that for every such basis element $v_{ij}^{\alpha}(x)$, there exists a skew-symmetric polynomial matrix
@@
$h(x)=B(x)x.$ (7)
These two facts (6) and (7) will be established in Proposition 3.15 and Lemma 3.17, respectively.
@@ 3.1 Proof for the Special Case: $x^{\top}g(x)=0$ for all $x\in\mathbb{R}^{n}$
$\mathbf{Q}_{[d]}^{n}:=\oplus_{m=1}^{d}\mathbf{Q}_{(m)}^{n}.$
@@
$\mathcal{V}:=\left\{v_{ij}^{\alpha}(x)=x^{\alpha+{e}_{j}}e_{i}-x^{\alpha+{e}_{i}}{e}_{j}\mid|\alpha|=d\right.\left.1\leq i<j\leq n\right\}.$ (10)
@@
$g^{\prime}_{2}(x)=\sum_{|\alpha|=d}\sum_{j>2}c^{\prime}_{\alpha,j}x^{\alpha+j},$
@@
The following result is then an immediate consequence of Lemma 3.5.
@@
The space $\mathbf{Q}^{n}_{(d+1)}$ admits a basis:
$\left\{v_{i,i+1}^{\alpha}=x^{\alpha+{e}_{i+1}}e_{i}-x^{\alpha+e_{i}}{e}_{i+1}\mid|\alpha|=d,i\in[n-1]\right\}$
and its dimension is given by
$\dim\mathbf{Q}^{n}_{(d+1)}=(n-1)\binom{n+d}{d}$.
@@
Proof 3.9 (Proof of Proposition 2).
Given $g\in\mathbf{Q}^{n}_{(d+1)}$, we write
@@
where the entries of $g_{(m)}\in\mathbf{P}^{n}_{(m)}$ contain only monomials of degree $m$. It suffices to show that for each $m=0,\ldots,d+1$, there exists a skew-symmetric polynomial matrix $A_{m}(x)$ of degree at most $d$ such that $g_{(m)}(x)=A_{m}(x)x$. Consider two cases:
@@
With all the results above, we present Algorithm 1 to illustrate the procedure of deriving the skew-symmetric matrix $A^{\prime}(x)$ from $g(x)\in\mathbf{Q}_{[d+1]}^{n}$.
@@
3 for $m\leftarrow 1$ to $d+1$ do
4   construct multi-index $\alpha$, basis $v_{ij}^{\alpha}(x)$, with $|\alpha|=m$
@@
7      compute $c_{\alpha,j}$ such that $g_{(m),i}(x)=\Bigl{(}g_{(m)}(x)-\sum_{|\alpha|=m}\sum_{j>i}c_{\alpha,j}\,v_{ij}^{\alpha}(x)\Bigr{)}e_{i}^{\top}=0$
8
@@
10
11     set $g_{(m)}(x)\leftarrow g_{(m)}(x)-\displaystyle\sum_{|\alpha|=m}\sum_{j>i}c_{\alpha,j}^{(i)}\,v_{ij}^{\alpha}(x)$
12
@@
14   set $A^{\prime}_{(m)}(x)\leftarrow\displaystyle\sum_{i,j}\sum_{|\alpha|=m}c_{\alpha,j}^{(i)}\,A_{ij}^{\alpha}(x)$
15
16Compute $A^{\prime}(x)=\sum_{m=1}^{d+1}A^{\prime}_{(m)}$

### v2

@@ Abstract
Replicator dynamics have been widely used in evolutionary game theory to model how strategy frequencies evolve over time in large populations. The so-called payoff matrix encodes the pairwise fitness that each strategy obtains when interacting with every other strategy, and it solely determines the replicator dynamics. If the payoff matrix is unknown, we show in this paper that it cannot be inferred from observed strategy frequencies alone — distinct payoff matrices can induce the same replicator dynamics. We thus look for a canonical representative of the payoff matrix in the equivalence class. The main result of the paper is to show that for every polynomial replicator dynamics (i.e., the vector field is a polynomial), there always exists a skew-symmetric, polynomial payoff matrix that can induce the given dynamics.
@@ 1 Introduction
The so-called replicator dynamics, originally introduced in [hauert2002replicator], have been widely used in evolutionary game theory [weibull1997evolutionary, hofbauer1998evolutionary] to describe how strategy frequencies change over time in a large population. Consider the basic setting where every agent in the population has $n$ (common) pure strategies at its disposal at any time. Denote by $x_{i}(t)$ the proportion of the agents taking the $i$th strategy at time $t$. Let $x(t):=(x_{1}(t),\ldots,x_{n}(t))$ be the system state. Note that $\sum_{i=1}^{n}x_{i}(t)=1$ and $x_{i}(t)\geq 0$ for all $i=1,\ldots,n$, so $x(t)$ is a probability vector. The agents in the population randomly pair with each other to play symmetric games whose payoff matrix is given by $H(x(t))\in\mathbb{R}^{n\times n}$ (i.e., the payoff matrix is a function of the current state $x(t)$). The payoff matrix encodes the competitive/cooperative nature between different strategies. Specifically, the $ij$th entry of $H(x(t))$ represents the payoff received by any agent using strategy $i$ when it interacts with an opponent using strategy $j$ with $x(t)$ the population state.
When playing the symmetric game, each agent evaluates the incremental benefit of abandoning its current strategy $i$ in favor of an alternative $k$. Conditional on facing an opponent who plays $j$, this benefit is the payoff difference $H_{kj}(x(t))-H_{ij}(x(t))$. Note that the strategy revision depends exclusively on such differences. Thus, it is invariant to any column-wise uniform shift of the payoff matrix (i.e., only the relative values of the entries of the $j$th column of $H(x(t))$ matter). Agents then follow a prescribed learning rule (e.g., pairwise proportional comparison and imitation driven by dissatisfaction [sandholm2010population]) to revise their strategies. Taking an appropriate limit (see, e.g., [kara2023differential] for details), we arrive at the following ordinary differential equation for $x(t)$:
@@
where $p(x)=(p_{1}(x),\ldots,p_{n}(x)):=H(x)x$ is known as the payoff function [park2019payoff] and $\bar{p}(x):=\sum_{j=1}^{n}x_{j}p_{j}(x)$ is the average payoff. The dynamics of $x(t)$ evolve on the standard simplex $\Delta^{n-1}:=\{x\in\mathbb{R}^{n}\mid x_{i}\geq 0\mbox{ and }\sum_{i=1}^{n}x_{i}=1\}$.
In addition to population games, we note that replicator dynamics also arise in the context of online learning [falniowski2025discrete, mertikopoulos2018cycles, sorin2020replicator] and reinforcement learning [hennes2020neural, paul2024multi]. To elaborate, consider the basic setting of online learning where a single agent has $n$ actions at its disposal, and its goal is to learn a mixed strategy $x(t)\in\Delta^{n-1}$ in real-time to minimize the expected regret, with the action-utility vector $p(x(t))\in\mathbb{R}^{n}$ revealed at each time-step. For certain learning rule such as FoReL [mertikopoulos2018cycles], it has been shown that the update equation of $x(t)$ can be approximated by the replicator dynamics as the continuous-time limit.
@@
In this paper, we take a different perspective: Assuming that one can only observe the dynamics of $x(t)$, we investigate whether it is feasible to identify the unknown payoff matrix $H:\Delta^{n-1}\to\mathbb{R}^{n\times n}$? Said in other words, we ask for any given replicator dynamics, is there a unique payoff matrix associated with it? If the answer is negative, then is there a “canonical” representative (we will make the canonicity clear)? The problems we pose play important roles in the setting where a monitor in their study of ecosystem aims to uncover the competitive/cooperative relations between different species, or, the setting where a central planner in the congestion game aims to understand the agents’ incentives against others’ choices of strategies. Specifically, we show that for a generalized class of replicator systems $\dot{x}=\operatorname{diag}(x)\,g(x),x\in\Delta^{n-1},$ where $g:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a polynomial vector field satisfying $x^{\!\top}g(x)=0$ for all $x\in\Delta^{n-1}$, there exists a skew-symmetric matrix $A(x)$ such that $g(x)=A(x)\,x,\text{with }A(x)=-A(x)^{\!\top}$. In other words, we show that every population game has its own zero-sum game representation. In particular, the skew-symmetric payoff matrices capture pure competition in which every gain for one strategy is matched by an equal loss to another. This structure yields a benchmark devoted to competitive dynamics while remaining tractable for the analysis of stability, coexistence [mai2018cycles], and learning behavior [mertikopoulos2018cycles]. For constant payoffs, the canonical case is the three-strategy Rock-Paper-Scissors (RPS) games and their $n$-strategy generalizations. The associated replicator dynamics exhibit recurrent interior orbits on the simplex and constitute the standard framework in which population permanence [amann1985permanence, hofbauer1987permanence] has been rigorously established.
@@
The replicator dynamics has been studied in various scenarios. In [mai2018cycles, skoulakis2021evolutionary], the authors investigated the co-evolution of the replicator system where the payoff matrix is generated from a dynamic system based on a generalized RPS game. The natural transformation [hofbauer1998evolutionary] was performed to study the cyclic behavior. In [park2019payoff, park2019population], the authors studied a more general setting where the payoff function is generated from a dynamic system, and the passivity theorem [pavel2022dissipativity, arcak2020dissipativity] is used to find the equilibria of the system. The most related work to ours is the static payoff model [park2018passivity], where the payoff is a static function. For $p(x)$ being static, the dynamic behavior of (1) has been extensively investigated in the literature. Many existing works have focused on the case where the payoff matrix $H(x)$ is constant: For $H$ a symmetric matrix, it has been shown [sandholm2010population] that the replicator dynamics (1) are gradient flows of the potential function $\bar{p}(x)$ with respect to the Shahshahani metric. For $H$ skew-symmetric, it has been shown [biggar2024attractor] that there is a unique global attractor of (1) (roughly speaking, an attractor is a minimal compact set that is forward invariant and asymptotically stable). For the case where $H$ is a 33-by-33 skew-symmetric matrix, a complete characterization of the attractors has been obtained by Zeeman [zeeman2006population]. In fact, more has been shown in [zeeman2006population] as well as in the follow-up work [bomze1983lotka, bomze1995lotka] for $H$ a general 33-by-33 matrix. Amongst others, they have presented, as a conjecture with partial proofs, all possible phase portraits (49 in total) for the resulting replicator dynamics with isolated equilibria and periodic orbits.
@@
The literature is relatively sparse for the case where the payoff matrix $H(x)$ depends on the state $x$, which is more realistic for a number of settings. For example, such dependence arises naturally in modeling the ecosystem where the interactions among different species are often in higher order [battiston2020networks] (in this context, the $x_{i}(t)$’s represent the normalized population densities of the species). Also, for congestion games [sandholm2010population] where different strategies represent different routes between the destinations, it is easy to see that the payoff matrix $H(x)$ depends on $x$ (the $x_{i}$’s represent the strategy frequencies). We further mention the demand-response program for smart grids [park2019payoff, Example 2] where the payoff matrix $H(x)$ depends on the user-strategy frequencies. The analysis of replicator dynamics for non-constant $H(x)$ is, however, much more involved and can quickly lose tractability. Results have been obtained for some special 33-by-33 payoff matrices $H(x)$ in [wesson2015replicator] where $H(x)$ is an affine function. Specifically, the authors have exhibited all the equilibria of the resulting replicator dynamics, analyzed their local stability by evaluating the Jacobian matrices, and carried out numerical studies for the phase portrait.
@@
We introduce the notation that will be used throughout the analysis. We define the hyperplane
$\mathbf{H}:=\{x\in\mathbb{R}^{n}|\sum_{i=1}^{n}x_{i}=1\}.$
Given arbitrary integer $d>0$, we let $\mathbf{\Omega}_{[d]}^{k\times k}$ be the vector space of all $k\times k$ skew-symmetric matrices $A$ with $A^{\top}=-A$, whose entries are polynomials in $x\in\mathbb{R}^{n}$ of degree at most $d$, and $\boldsymbol{\Omega}_{(d)}^{k\times k}$ denote the homogeneous subspace of degree $d$. If $d=0$, we write $\mathbf{\Omega}^{n\times n}$ for simplicity.
@@
${\mathbf{\Omega}^{k\times k}_{[d]}=\oplus_{m=0}^{d}\mathbf{\Omega}^{k\times k}_{(m)}\ \mbox{ and }\ }\mathbf{P}^{k\times\ell}_{[d]}=\oplus_{m=0}^{d}\mathbf{P}^{k\times\ell}_{(m)}.$
For $d<0$, we let the corresponding vector space be the trivial space, i.e., the space with only the 0 vector.
Denote by $\mathbb{N}_{0}$ the set of nonnegative integers. We use multi-index $x^{\alpha}$ to denote the monomial $x^{\alpha}:=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}$, where $\alpha=\left(\alpha_{1},\ldots,\alpha_{n}\right)\in\mathbb{N}^{n}_{0}$ and $|\alpha|:=\sum_{i=1}^{n}\alpha_{i}$ is the degree of the monomial. We say that a scalar polynomial contains a monomial if the coefficient of the monomial in the standard multivariate expansion of the polynomial is non-zero. The degree of the polynomial is by convention the largest degree of a monomial contained in it. For a vector- or matrix-valued polynomial, we define its degree to be the largest degree among all of its entries.
2 Motivating Example and Main Result
@@ 2.1 Motivating Example
$\phi:H(x)\mapsto f(x):=\operatorname{diag}(x)(H(x)x-x^{\top}H(x)x\mathbf{1}),$ (2)
@@ 2.2 Main Result
then, there exists a polynomial map $A:\mathbb{R}^{n}\to\mathbf{\Omega}^{n\times n}$ of degree at most $d$ such that $g(x)=A(x)x$, for all $x\in\mathbf{H}$.
@@
To wit, $x^{\top}g(x)$ is a polynomial function (and hence, analytic) and it is identically zero on $\Delta^{n-1}$, whose interior is open in $\mathbf{H}$. It then follows from the identity theorem [scheidemann2005introduction] that $x^{\top}g(x)=0$ for all $x\in\mathbf{H}$.
@@
It is clear that all such $g(x)$ of degree $(d+1)$ form a vector space, which we denote by $\mathbf{Q}_{(d+1)}^{n}$. We will provide a complete characterization of the vector space by exhibiting a spanning set of it. In particular, we will show in Lemma 3.5 that each element takes the following form:
@@
where $\alpha$ is some multi-index with $|\alpha|=d$. It then follows that for every such element $v_{ij}^{\alpha}(x)$, there exists a skew-symmetric polynomial matrix
@@
$h(x)=B(x)x,{\text{ for all }x\in\mathbf{H}.}$ (7)
These two facts (6) and (7) will be established in Lemma 3.16 and Lemma 3.18, respectively.
@@ 3.1 Proof for the Special Case: $x^{\top}g(x)=0$ for all $x\in\mathbb{R}^{n}$
$\mathbf{Q}_{[d]}^{n}:=\oplus_{m=0}^{d}\mathbf{Q}_{(m)}^{n}.$
@@
$\mathcal{V}:=\left\{v_{ij}^{\alpha}(x)=x^{\alpha+{e}_{j}}e_{i}-x^{\alpha+{e}_{i}}{e}_{j}\mid|\alpha|=d,\right.\left.1\leq i<j\leq n\right\}.$ (10)
@@
$g^{\prime}_{2}(x)=\sum_{|\alpha|=d}\sum_{j>2}c^{\prime}_{\alpha,j}x^{\alpha+e_{j}},$
@@
The following result computes the dimension of $\mathbf{Q}^{n}_{(d+1)}$.
@@
The dimension of $\mathbf{Q}^{n}_{(d+1)}$ is given by
$\dim\mathbf{Q}_{(d+1)}^{n}=n\binom{n+d}{d+1}-\binom{n+d+1}{d+2}$.
Proof 3.8.
Consider the linear map $L:\mathbf{P}_{(d+1)}^{n}\rightarrow\mathbf{P}_{(d+2)}^{1}$ given by $L(g(x)):=x^{\top}g(x)$. It follows that $\mathbf{Q}_{(d+1)}^{n}$ is the kernel of $L$. We show that $L$ is surjective. For any given $q(x)\in\mathbf{P}_{(d+2)}^{1}$, let $g(x):=\frac{1}{d+2}\frac{\partial q(x)}{\partial x}\in\mathbf{P}_{(d+1)}^{n}$. Since $q(x)$ is a homogeneous polynomial of degree $(d+2)$, it follows from Euler’s identity [euler2012elements] that $L(g(x))=q(x)$. Then, by the rank-nullity theorem, we have that
$\dim\mathbf{Q}_{(d+1)}^{n}=\dim\mathbf{P}_{(d+1)}^{n}-\dim\mathbf{P}_{(d+2)}^{1}.$ (13)
To evaluate $\dim\mathbf{P}_{(d+1)}^{n}$, we note that the number of monomials $x^{\alpha}$, with $|\alpha|=d+1$, is $\binom{n+d}{d+1}$ (known as the stars-and-bars problem [flajolet2009analytic]), so
$\dim\mathbf{P}_{(d+1)}^{n}=n\binom{n+d}{d+1}$. (14)
Similarly,
$\dim\mathbf{P}_{(d+2)}^{1}=\binom{n+d+1}{d+2}$. (15)
The corollary then follows from (13), (14) and (15). ∎
@@
Proof 3.10 (Proof of Proposition 2).
Given $g\in\mathbf{Q}^{n}_{[d+1]}$, we write
@@
where the entries of $g_{(m)}\in\mathbf{P}^{n}_{(m)}$ contain only monomials of degree $m$. It suffices to show that for each $m=0,\ldots,d+1$, there exists a skew-symmetric polynomial matrix $A_{m}(x)$ of degree at most (m1)(m-1) such that $g_{(m)}(x)=A_{m}(x)x$. Consider two cases:
@@
With all the results above, we present Algorithm 1 to illustrate the procedure of deriving the skew-symmetric matrix $A^{\prime}(x)$ from $g(x)\in\mathbf{Q}_{[d+1]}^{n}$. In the for-loop (line 3103-10), computing the coefficients reduces to solving linear systems whose number of equalities matches the number of basis vectors (via elimination). By Corollary 3.7, the dimension of $\mathbf{Q}_{[d+1]}^{n}$ is given by
$N:=\dim\mathbf{Q}_{[d+1]}^{n}=\sum_{k=0}^{d}\dim\mathbf{Q}_{(k+1)}^{n}=n\sum_{k=0}^{d}\binom{n+k}{k+1}-\sum_{k=0}^{d}\binom{n+k+1}{k+2}=(n-1)\binom{n+d+1}{d+1}-\binom{n+d+1}{d+2}+1.$
Using Gaussian elimination, solving such systems requires $O(N^{3})$ arithmetic operations [golub2013matrix]. Hence, Algorithm 1 has overall complexity of $O(N^{3})$, which is polynomial in $n$ and exponential in $d$.
@@
3
4for $m\leftarrow 1$ to $d+1$ do
5   construct multi-index $\alpha$, spanning set of $v_{ij}^{\alpha}(x)$, with $|\alpha|=m-1$
@@
8      compute $c_{\alpha,j}$ such that $e_{i}^{\top}\Bigl(g_{(m)}(x)-\sum_{|\alpha|=m-1}\sum_{j>i}c_{\alpha,j}\,v_{ij}^{\alpha}(x)\Bigr)=0$
9
@@
11
12     set $g_{(m)}(x)\leftarrow g_{(m)}(x)-\displaystyle\sum_{|\alpha|=m-1}\sum_{j>i}c_{\alpha,j}^{(i)}\,v_{ij}^{\alpha}(x)$
13
@@
15   set $A^{\prime}_{(m)}(x)\leftarrow\displaystyle\sum_{i,j}\sum_{|\alpha|=m-1}c_{\alpha,j}^{(i)}\,A_{ij}^{\alpha}(x)$
16
17Compute $A^{\prime}(x)=\sum_{m=1}^{d+1}A^{\prime}_{(m)}$